OpenStudy (anonymous):

2x=14 What can I do to isolat the variable? Multiply or Divide?

OpenStudy (aravindg):

div by 2

OpenStudy (aravindg):

:) medal plzz

OpenStudy (anonymous):

2x / 2 = 14 / 2 x = 7

OpenStudy (amistre64):